### v4

@@ Appendix A Auxiliary Results
$=\frac{h_{L}^{-\frac{9}{4}}}{(1-\alpha)^{2}}\frac{h_{\ell}}{N_{\ell}}h_{\ell}^{-1}\mathbb{V}\mathrm{ar}\big((X_{h_{\ell}}-\xi^{h_{\ell}}_{\star})^{+}-(X_{h_{\ell-1}}-\xi^{h_{\ell-1}}_{\star})^{+}\big).$
Via (27) , (90) and Cesàro’s lemma (‘$0/0$’ version),
@@
$\quad\times\lim_{L\uparrow\infty}\bigg(\sum_{\ell=1}^{L}\frac{h_{\ell}}{N_{\ell}}\bigg)^{-1}\bigg(\sum_{\ell=1}^{L}\frac{h_{\ell}}{N_{\ell}}h_{\ell}^{-1}\mathbb{V}\mathrm{ar}\big((X_{h_{\ell}}-\xi^{h_{\ell}}_{\star})^{+}-(X_{h_{\ell-1}}-\xi^{h_{\ell-1}}_{\star})^{+}\big)\bigg)$
@@
$\quad\times\lim_{L\uparrow\infty}\bigg(\sum_{\ell=1}^{L}h_{\ell}^{\frac{1}{4}}\bigg)^{-1}\bigg(\sum_{\ell=1}^{L}h_{\ell}^{\frac{1}{4}}h_{\ell}^{-1}\mathbb{V}\mathrm{ar}\big((X_{h_{\ell}}-\xi^{h_{\ell}}_{\star})^{+}-(X_{h_{\ell-1}}-\xi^{h_{\ell-1}}_{\star})^{+}\big)\bigg)$
$=\frac{h_{0}^{\frac{1}{4}}}{(1-\alpha)^{2}M^{\frac{1}{4}}}\mathbb{V}\mathrm{ar}(\mathds{1}_{X_{0}>\xi^{0}_{\star}}\,G).$
@@
$S^{1,2}_{L}=S^{2,1}_{L}\mathrel{{\mathop{\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow}\limits^{\mathbb{P}\text{\rm-as}}}}0\quad\mbox{ as }\quad L\uparrow\infty.$
@@ Appendix G Square Convergence of the Averaged MLSA Algorithm
To fully apprehend its properties, we hereby study the $L^{2}(\mathbb{P})$-error and complexity of the VaR AMLSA algorithm (26).
@@ G.1 Convergence Rate Analysis
    Let $\big\{G_{\ell}\coloneqq h_{\ell}^{-\frac{1}{2}}(X_{h_{\ell}}-X_{h_{\ell-1}}),\ell\geq 1\big\}$, define $\{F_{X_{h_{\ell-1}}\,|\,G_{\ell}=g}\colon x\mapstochar\rightarrow\mathbb{P}(X_{h_{\ell-1}}\leq x\,|\,G_{\ell}=g),g\in\mathrm{supp}(\mathbb{P}_{G_{\ell}}),\ell\geq 1\}$ and consider the random variables $\{K_{\ell}\coloneqq K_{\ell}(G_{\ell}),\ell\geq 1\}$, where
    $K_{\ell}(g)\coloneqq\sup_{x\neq y}\,\frac{\big|F_{X_{h_{\ell-1}}\,|\,G_{\ell}=g}(x)-F_{X_{h_{\ell-1}}\,|\,G_{\ell}=g}(y)\big|}{|x-y|},\quad\ell\geq 1,g\in\mathrm{supp}(\mathbb{P}_{G_{\ell}}).$
@@
$\mathbb{E}\big[\big(\overline{\xi}^{\text{\tiny\rm ML}}_{\mathbf{N}}$ $-\xi^{h_{L}}_{\star}\big)^{2}\big]\leq C\bigg(\gamma_{N_{0}}+\bigg(\sum_{\ell=1}^{L}\frac{1}{N_{\ell}\sqrt{\gamma_{N_{\ell}}}}\bigg)^{2}$ (95)
@@
We come back to the decomposition (94) and analyze each term separately.
@@
$|V_{0}^{\prime\prime}(\xi^{0}_{\star})-V_{h_{\ell}}^{\prime\prime}(\xi^{h_{\ell}}_{\star})|\leq C\big(h_{\ell}^{(\frac{1}{4}+\delta_{0})\wedge 1}+h_{\ell}\big).$ (96)
Using (36), (96) and Lemma A.3(ii),
@@
From (37), (47), the uniform Lipschitz regularity of $\{V_{h_{\ell}}^{\prime\prime},\ell\geq 1\}$, stemming from Assumption 1.1(iv), and Lemma G.1(i), we get
@@
Recall that $\big\{\frac{1}{N_{\ell}}\sum_{k=1}^{N_{\ell}}(\rho^{h_{\ell}}_{k}-\rho^{h_{\ell-1}}_{k}),\ell\geq 1\big\}$ are independent and centered and that, at each level $\ell\geq 1$, $\{\rho^{h_{\ell}}_{k}-\rho^{h_{\ell-1}}_{k},k\geq 1\}$ are $\{\mathcal{F}^{h_{\ell}}_{k},k\geq 1\}$-martingale increments. Using (49) and a comparison between series and integrals,
@@
Since $\big\{\frac{1}{N_{\ell}}\sum_{k=1}^{N_{\ell}}(e^{h_{\ell}}_{k}-e^{h_{\ell-1}}_{k}),\ell\geq 1\big\}$ are independent and centered and, for any level $\ell\geq 1$, $\big\{e^{h_{\ell}}_{k}-e^{h_{\ell-1}}_{k},k\geq 1\big\}$ are $\{\mathcal{F}^{h_{\ell}}_{k},k\geq 1\}$-martingale increments, using (91) and Lemma G.1(ii),
@@ G.2 Complexity Analysis
The global error of the VaR AMLSA scheme (26) decomposes into a statistical and a bias errors:
@@
for some positive constant $C<\infty$ independent of $L$. To retrieve the optimal iterations amounts $\mathbf{N}=\{N_{\ell},0\leq\ell\leq L\}$, we test out several leading term candidates for the upper bound (95), of which we retain three compelling cases. We minimize the computational cost while constraining the chosen candidate to an order of $\varepsilon^{2}$.
@@
$\begin{array}[]{lll}\operatorname*{minimize}_{N_{0},\dots,N_{L}>0}&\sum_{\ell=0}^{L}N_{\ell}h_{\ell}^{-1}\\ \operatorname{subject\penalty 10000\ to}&\sum_{\ell=0}^{L}h_{\ell}^{(1+4\delta_{0}\wedge 3)/4}\sqrt{\gamma_{N_{\ell}}}=C^{-1}\varepsilon,\end{array}$
@@
to guarantee that the upper estimate (95) is of order $\varepsilon^{2}$. The corresponding complexity satisfies
$\operatorname{Cost}^{\beta}_{\text{\rm\tiny AMLSA}}\leq C\varepsilon^{-\frac{2}{\beta}}\mathrel{{\mathop{\rightarrow}\limits}}\varepsilon^{-3}\quad\mbox{ as }\quad\beta\uparrow\frac{2}{3}.$
@@
$\begin{array}[]{lll}\operatorname*{minimize}_{N_{0},\dots,N_{L}>0}&\sum_{\ell=0}^{L}N_{\ell}h_{\ell}^{-1}\\ \operatorname{subject\penalty 10000\ to}&\sum_{\ell=0}^{L}N_{\ell}^{-1}\gamma_{N_{\ell}}^{-\frac{1}{2}}=C^{-1}\varepsilon.\end{array}$
@@
are met, one attains the order $\varepsilon^{2}$ for the global $L^{2}(\mathbb{P})$-error (95), with an optimal cost satisfying
$\operatorname{Cost}^{\beta}_{\text{\rm\tiny AMLSA}}\leq C\varepsilon^{-\frac{2}{2-\beta}-1}\mathrel{{\mathop{\rightarrow}\limits}}\varepsilon^{-\frac{5}{2}}\quad\mbox{ as }\quad\beta\downarrow\frac{2}{3}.$
@@
$\begin{array}[]{lll}\operatorname*{minimize}_{N_{0},\dots,N_{L}>0}&\sum_{\ell=0}^{L}N_{\ell}h_{\ell}^{-1}\\ \operatorname{subject\penalty 10000\ to}&\sum_{\ell=0}^{L}h_{\ell}^{\frac{1}{2}}N_{\ell}^{-1}=C^{-1}\varepsilon^{2}.\end{array}$
@@
The convergence analysis of this last heuristic is deferred to Theorem 4.1 and the ensuing complexity is discussed in Theorem $4.$

### v5

@@ Appendix A Auxiliary Results
$=\frac{h_{L}^{-\frac{9}{4}}}{(1-\alpha)^{2}}\frac{h_{\ell}}{N_{\ell}}h_{\ell}^{-1}\operatorname{\mathbb{V}ar}\big((X_{h_{\ell}}-\xi^{h_{\ell}}_{\star})^{+}-(X_{h_{\ell-1}}-\xi^{h_{\ell-1}}_{\star})^{+}\big).$
Via (4.2) , (E.14) and Cesàro’s lemma (‘$0/0$’ version),
@@
$\quad\times\lim_{L\uparrow\infty}\bigg(\sum_{\ell=1}^{L}\frac{h_{\ell}}{N_{\ell}}\bigg)^{-1}\bigg(\sum_{\ell=1}^{L}\frac{h_{\ell}}{N_{\ell}}h_{\ell}^{-1}\operatorname{\mathbb{V}ar}\big((X_{h_{\ell}}-\xi^{h_{\ell}}_{\star})^{+}-(X_{h_{\ell-1}}-\xi^{h_{\ell-1}}_{\star})^{+}\big)\bigg)$
@@
$\quad\times\lim_{L\uparrow\infty}\bigg(\sum_{\ell=1}^{L}h_{\ell}^{\frac{1}{4}}\bigg)^{-1}\bigg(\sum_{\ell=1}^{L}h_{\ell}^{\frac{1}{4}}h_{\ell}^{-1}\operatorname{\mathbb{V}ar}\big((X_{h_{\ell}}-\xi^{h_{\ell}}_{\star})^{+}-(X_{h_{\ell-1}}-\xi^{h_{\ell-1}}_{\star})^{+}\big)\bigg)$
$=\frac{h_{0}^{\frac{1}{4}}}{(1-\alpha)^{2}M^{\frac{1}{4}}}\operatorname{\mathbb{V}ar}(\mathds{1}_{X_{0}>\xi^{0}_{\star}}\,G).$
@@
$S^{1,2}_{L}=S^{2,1}_{L}\mathrel{{\mathop{\longrightarrow}\limits^{\operatorname{\mathds{P}\text{-as}}}}}0\quad\mbox{ as }\quad L\uparrow\infty.$
@@ Appendix G Square Convergence of the Averaged MLSA Algorithm
To fully apprehend its properties, we hereby study the $L^{2}(\mathbb{P})$-error and complexity of the VaR AMLSA algorithm (4.1).
@@ G.1 Convergence Rate Analysis
    Let $\big\{G_{\ell}\coloneqq h_{\ell}^{-\frac{1}{2}}(X_{h_{\ell}}-X_{h_{\ell-1}}),\ell\geq 1\big\}$, define $\{F_{X_{h_{\ell-1}}\,|\,G_{\ell}=g}\colon x\mapsto\mathbb{P}(X_{h_{\ell-1}}\leq x\,|\,G_{\ell}=g),g\in\operatorname*{supp}(\mathbb{P}_{G_{\ell}}),\ell\geq 1\}$ and consider the random variables $\{K_{\ell}\coloneqq K_{\ell}(G_{\ell}),\ell\geq 1\}$, where
    $K_{\ell}(g)\coloneqq\sup_{x\neq y}\,\frac{\big|F_{X_{h_{\ell-1}}\,|\,G_{\ell}=g}(x)-F_{X_{h_{\ell-1}}\,|\,G_{\ell}=g}(y)\big|}{|x-y|},\quad\ell\geq 1,g\in\operatorname*{supp}(\mathbb{P}_{G_{\ell}}).$
@@
$\mathbb{E}\big[\big(\overline{\xi}^{\text{\tiny\rm ML}}_{\mathbf{N}}$ $-\xi^{h_{L}}_{\star}\big)^{2}\big]\leq C\bigg(\gamma_{N_{0}}+\bigg(\sum_{\ell=1}^{L}\frac{1}{N_{\ell}\sqrt{\gamma_{N_{\ell}}}}\bigg)^{2}$ (G.1)
@@
We come back to the decomposition (F.2) and analyze each term separately.
@@
$|V_{0}^{\prime\prime}(\xi^{0}_{\star})-V_{h_{\ell}}^{\prime\prime}(\xi^{h_{\ell}}_{\star})|\leq C\big(h_{\ell}^{(\frac{1}{4}+\delta_{0})\wedge 1}+h_{\ell}\big).$ (G.2)
Using (B.2), (G.2) and Lemma A.3(ii),
@@
From (B.3), (B.13), the uniform Lipschitz regularity of $\{V_{h_{\ell}}^{\prime\prime},\ell\geq 1\}$, stemming from Assumption 1.1(iv), and Lemma G.1(i), we get
@@
Recall that $\big\{\frac{1}{N_{\ell}}\sum_{k=1}^{N_{\ell}}(\rho^{h_{\ell}}_{k}-\rho^{h_{\ell-1}}_{k}),\ell\geq 1\big\}$ are independent and centered and that, at each level $\ell\geq 1$, $\{\rho^{h_{\ell}}_{k}-\rho^{h_{\ell-1}}_{k},k\geq 1\}$ are $\{\mathcal{F}^{h_{\ell}}_{k},k\geq 1\}$-martingale increments. Using (B.15) and a comparison between series and integrals,
@@
Since $\big\{\frac{1}{N_{\ell}}\sum_{k=1}^{N_{\ell}}(e^{h_{\ell}}_{k}-e^{h_{\ell-1}}_{k}),\ell\geq 1\big\}$ are independent and centered and, for any level $\ell\geq 1$, $\big\{e^{h_{\ell}}_{k}-e^{h_{\ell-1}}_{k},k\geq 1\big\}$ are $\{\mathcal{F}^{h_{\ell}}_{k},k\geq 1\}$-martingale increments, using (E.15) and Lemma G.1(ii),
@@ G.2 Complexity Analysis
The global error of the VaR AMLSA scheme (4.1) decomposes into a statistical and a bias errors:
@@
for some positive constant $C<\infty$ independent of $L$. To retrieve the optimal iterations amounts $\mathbf{N}=\{N_{\ell},0\leq\ell\leq L\}$, we test out several leading term candidates for the upper bound (G.1), of which we retain three compelling cases. We minimize the computational cost while constraining the chosen candidate to an order of $\varepsilon^{2}$.
@@
$\begin{array}[]{lll}\operatorname{minimise}_{N_{0},\dots,N_{L}>0}&\sum_{\ell=0}^{L}N_{\ell}h_{\ell}^{-1}\\ \operatorname{subject\penalty 10000\ to}&\sum_{\ell=0}^{L}h_{\ell}^{(1+4\delta_{0}\wedge 3)/4}\sqrt{\gamma_{N_{\ell}}}=C^{-1}\varepsilon,\end{array}$
@@
to guarantee that the upper estimate (G.1) is of order $\varepsilon^{2}$. The corresponding complexity satisfies
$\operatorname{Cost}^{\beta}_{\text{\rm\tiny AMLSA}}\leq C\varepsilon^{-\frac{2}{\beta}}\mathrel{{\mathop{\to}\limits}}\varepsilon^{-3}\quad\mbox{ as }\quad\beta\uparrow\frac{2}{3}.$
@@
$\begin{array}[]{lll}\operatorname{minimise}_{N_{0},\dots,N_{L}>0}&\sum_{\ell=0}^{L}N_{\ell}h_{\ell}^{-1}\\ \operatorname{subject\penalty 10000\ to}&\sum_{\ell=0}^{L}N_{\ell}^{-1}\gamma_{N_{\ell}}^{-\frac{1}{2}}=C^{-1}\varepsilon.\end{array}$
@@
are met, one attains the order $\varepsilon^{2}$ for the global $L^{2}(\mathbb{P})$-error (G.1), with an optimal cost satisfying
$\operatorname{Cost}^{\beta}_{\text{\rm\tiny AMLSA}}\leq C\varepsilon^{-\frac{2}{2-\beta}-1}\mathrel{{\mathop{\to}\limits}}\varepsilon^{-\frac{5}{2}}\quad\mbox{ as }\quad\beta\downarrow\frac{2}{3}.$
@@
$\begin{array}[]{lll}\operatorname{minimise}_{N_{0},\dots,N_{L}>0}&\sum_{\ell=0}^{L}N_{\ell}h_{\ell}^{-1}\\ \operatorname{subject\penalty 10000\ to}&\sum_{\ell=0}^{L}h_{\ell}^{\frac{1}{2}}N_{\ell}^{-1}=C^{-1}\varepsilon^{2}.\end{array}$
@@
The convergence analysis of this last heuristic is deferred to Theorem 4.1 and the ensuing complexity is discussed in Theorem 4.3.